OpenStudy (sneakysquash):

Please graph- y=1/3x+2

OpenStudy (codysek98):

Do you know what a y-intercept is and where you can find the y-intercept in your equation?

OpenStudy (sneakysquash):

Yeah I do, the y intercept would be1/3x yeah?

OpenStudy (codysek98):

Other way around. The y-intercept is what number you get when you plug 0 in for x.

OpenStudy (codysek98):

So your y-intercept would be the 2 at the end there. Now for your slope, what do you think the slope is?

OpenStudy (sneakysquash):

I honestly don't know man, could you please graph it out for me?

OpenStudy (codysek98):

I'll explain the parts of the equation for you and what they are. I can't just graph it for you, it'll be giving you the answer without you learning it.

OpenStudy (studygurl14):

@sneakysquash you can easily graph it at www.desmos.com/calculator

OpenStudy (studygurl14):

But you should listen to @codysek98 so you can learn how to do it

OpenStudy (sneakysquash):

Okay then.

OpenStudy (will.h):

\[\frac{ 1 }{ 3 }x + 2\] or \[\frac{ 1 }{ 3x } + 2\]

OpenStudy (studygurl14):

the first one @Will.H

OpenStudy (will.h):

oh then create 2 point when x = 0 and the other when y = 0 and that's all

OpenStudy (codysek98):

So in your equation, y=1/3x+2. 1/3 = m which is your slope. 2 = y-intercept To graph the equation, you want to start with your y-intercept, which is (0,2). To graph more points (usually 3 or 4 points is good unless told otherwise), you want to use your slope. The slope is written in rise/run, so you would want to go up 1 and right 3 from your y-intercept and you can also go down 1 and left 3.

OpenStudy (codysek98):

I'm very bad at wording things but that's basically the process^

OpenStudy (sneakysquash):

Okay, thanks dude.